# DINO-R1: Incentivizing Reasoning Capability in Vision Foundation Models

### v1

@@ 1 Introduction
Recent advancements in large reasoning models (LRMs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10], exemplified by the impressive performance of DeepSeek-R1 [8, 1], have demonstrated remarkable capabilities across complex reasoning tasks like math reasoning and coding. This breakthrough is largely driven by innovative reinforcement (RL) learning strategies such as Group Relative Policy Optimization (GRPO) [1]. By iteratively generating synthetic data and optimizing reasoning models through verifiable rewards, DeepSeek-R1 has attained superior reasoning abilities competitive with state-of-the-art proprietary models, such as OpenAI o1, significantly reshaping the language modeling landscape. Despite these impressive advancements, however, the equivalent advancement of reasoning remains notably limited in vision foundation models (VFMs) [11, 12, 13, 14, 15, 16, 17]. Current VFMs largely rely on supervised training paradigms focused on predefined visual categories [18, 19, 20] or self-supervised objectives [21, 22, 23]. These conventional supervised methods inherently lack robust reasoning mechanisms, limiting their ability to effectively generalize to novel, ambiguous, or high-variance scenarios frequently encountered in practical, real-world applications.
@@
One emerging and increasingly important scenario in VFMs is visual prompting [24, 16, 25, 15, 26], a new paradigm wherein users specify detection targets using visual exemplars. This approach offers substantial practical utility in broad applications such as auto-labeling, industrial inspection, and robotic manipulation. [27, 28, 29, 30, 31, 32, 33, 34, 35, 36] However, training visual-prompting models poses new challenges despite their practical relevance due to the high diversity and intra-class variation among visual exemplars. Compared to their language-prompted counterparts [14, 37, 38, 26, 39], training recipes for visual prompting detectors remain largely underdeveloped. We have observed that training with supervised fine-tuning (SFT) alone typically struggles under these conditions, exhibiting unstable convergence, limited generalization to out-of-domain data, as well as poor alignment of query predictions with visual prompts (Sec. 4). These findings suggest that vanilla SFT is insufficient for effectively training visual prompting detectors, motivating us to explore fundamentally new training strategies capable of effectively reasoning over diverse visual inputs for robust generalization.
@@
Inspired by the recent breakthroughs in RL-based training frameworks for LRMs [6, 7, 8, 9, 1, 40, 41, 42, 2, 10, 3, 4, 5], which efficaciously exploit large-scale noisy training data, we aim to similarly unlock the potential of reasoning capabilities within pure vision models, e.g., VFMs. Yet, naïve application of language-based RL methods such as GRPO to vision presents non-trivial challenges. For one thing, GRPO assumes the model behaves as a probabilistic generator, explicitly sampling diverse outputs from learned distributions per input. In contrast, vision models typically produce deterministic structured predictions, making it nontrivial to optimize over a sampled output space. For another, GRPO’s KL-regularization that stabilizes training via constraining token-level output distributions in language models can not be easily translated to structured visual predictions due to fundamental differences in language and vision formulation.
@@ 2 Related Work
Vision foundation models (VFMs) [43, 16, 15, 44, 14, 37, 45, 46, 11, 12, 47, 24, 25, 48, 49] have enabled significant progress across a wide spectrum of visual tasks by learning general-purpose image representations from large-scale datasets. Among them, DETR [50] and its derivatives [51, 52, 13, 16, 14, 53, 15, 37, 38] formulate object detection as a set prediction problem using transformer-based architectures, offering open-set capabilities and strong performance on dense prediction tasks. Grounding DINO [14, 38], in particular, extends DETR to open-vocabulary detection by incorporating vision-language alignment, enabling phrase-grounded object detection via language prompts.
@@
Open-vocabulary detection [54](OVD) aims to recognize categories beyond the supervised training set by leveraging external knowledge sources such as pretrained text encoders or image-text pairs. Most existing OVD methods [55, 39, 53, 54, 56, 14] focus on language prompts to bridge the category gap, leaving the space of visual prompting—using visual exemplars instead of text—largely underexplored. Recent works [26, 16, 24, 25, 15] have explored using reference images or bounding boxes to ground object-level semantics. However, these methods often rely on inference-time conditioning and lack a robust training paradigm for learning from high-variance visual prompts.
@@
Reinforcement learning [41, 40, 42, 1, 6] has played a central role in fine-tuning large language models (LLMs)[2, 3, 57, 4, 5, 6, 8] via methods such as Reinforcement Learning from Human Feedback (RLHF) [41] and Group Relative Policy Optimization (GRPO)[1, 8]. These approaches enable models to align better with diverse, weakly-supervised, or ambiguous objectives. However, the application of reinforcement-style training to vision foundation models, especially for dense prediction tasks like object detection, remains underexplored. Our work bridges this gap by adapting GRPO principles to query-level learning in transformer-based object detectors through our proposed Group Relative Query Optimization.
@@ 3 Methodology
While language-prompted object detection has received growing attention in the vision-language community, the training strategy for visual prompt-based detection remains underexplored. To address the challenge of high-variance visual prompts and unlock the potential of prompt-guided detector, in this work, we introduce a novel training paradigm for the visual prompting object detection, Group Relative Query Optimization (GRQO). Built on top of the Grounding-DINO (G-DINO) [14] framework (§3.1), our approach integrates a visual prompt encoder (§3.2) and the GRQO mechanism (§3.3) to enhance query learning and improve detection robustness.
@@ 3.2 VIS-G-DINO
Following G-DINO’s architecture, we fuse image features and visual prompts via a multimodal feature enhancer $\mathcal{F}_{I,v}$ to obtain refined image features $\mathbf{I}^{\prime}$ and visual prompt features $\bar{\mathbf{v}}^{\prime}$. To guide detection process, we introduce a visual-guided query selection mechanism. Given refined image tokens $\mathbf{I}^{\prime}$ and visual prompt features $\bar{\mathbf{v}}^{\prime}$, we compute an image-prompt similarity matrix via dot product. For each image token, we take the maximum similarity across prompt axis as its objectness score, representing the likelihood that a prompted object exists at that location. We select the top-$N_{q}$ image tokens with the highest objectness scores as positional embeddings for the decoder queries. The top-$N_{q}$ indices selection can be expressed as:
@@ 3.3 Group Relative Query Optimization
Visual prompting detection demands that object queries align with highly diverse visual exemplars that share the same semantics. This setting introduces greater intra-class variance than language prompts, requiring the model to both memorize diverse appearances and generalize across unseen variations. Inspired by the generalization ability of GRPO [1] in LLMs community, we propose Group Relative Query Optimization (GRQO, Fig.2)—a novel training paradigm that enhances query quality and learning stability through group-based reward modeling and distributional regularization.
@@
where $\alpha$ and $\beta$ are scalar weights that balance the reward signal and regularization strength. GRQO incentivizes both query-level learning and stable objectness modeling.
@@ 4.1 Experiment Setup
We compare GRQO with standard supervised fine-tuning (SFT). All experiments are conducted using the MM-Grounding-DINO [38] implementation, which we adapt to support visual prompting. We use visual exemplars (images with user defined bounding boxes) as input prompts to guide the detection in the target image.
@@
We evaluate our approach under two settings: ❶ Zero-Shot. (Out-domain Evaluation). We conduct open-vocabulary detection where we train on Objects365 (O365) [20] and test on COCO [19], LVIS-minival [58], ODinW13 [45], and ODinW35 [45]. We train the model for 6 epochs as SFT baselines. For GRQO, we first train 1 epoch with SFT to obtain the reference model, then apply GRQO for 5 additional epochs. ❷ Fine-Tune. (In-domain Evaluation). We fine-tune on the COCO training split for 12 epochs and evaluate on the COCO validation set. For GRQO, we apply both pretrained weights of SFT and GRQO for evaluation.
@@ 4.2 Main Results
GDINO-T [14] O365,G,C 30 48.4 28.8 18.8 24.2 34.7 51.4 22.7 58.1
MM-GDINO-T [38] O365,G 30 50.4 35.7 28.1 30.2 42.0 45.3 20.2 58.2
MM-GDINO-L [38] O365v2,O,G 42 53.0 - - - - 75.3 35.2 -

### v2

@@ 1 Introduction
Recent advancements in large reasoning models (LRMs) grpo ; llama2 ; llama3 ; qwen ; qwen2 ; bi2024deepseek ; lu2024deepseek ; guo2025deepseek ; liu2024deepseek ; llava , exemplified by the impressive performance of DeepSeek-R1 guo2025deepseek ; grpo , have demonstrated remarkable capabilities across complex reasoning tasks like math reasoning and coding. This breakthrough is largely driven by innovative reinforcement (RL) learning strategies such as Group Relative Policy Optimization (GRPO) grpo . By iteratively generating synthetic data and optimizing reasoning models through verifiable rewards, DeepSeek-R1 has attained superior reasoning abilities competitive with state-of-the-art proprietary models, such as OpenAI o1, significantly reshaping the language modeling landscape. Despite these impressive advancements, however, the equivalent advancement of reasoning remains notably limited in vision foundation models (VFMs) sam ; sam2 ; dino ; gdino ; dinox ; dinov ; vit . Current VFMs largely rely on supervised training paradigms focused on predefined visual categories imagenet ; cocods ; o365 or self-supervised objectives dinov2 ; SimCLR ; G-SimCLR . These conventional supervised methods inherently lack robust reasoning mechanisms, limiting their ability to effectively generalize to novel, ambiguous, or high-variance scenarios frequently encountered in practical, real-world applications.
@@
One emerging and increasingly important scenario in VFMs is visual prompting trex ; dinov ; trex2 ; dinox ; cp-detr , a new paradigm wherein users specify detection targets using visual exemplars. This approach offers substantial practical utility in broad applications such as auto-labeling, industrial inspection, and robotic manipulation. app-visual-autovp ; app-visual-moka ; app-visual-multimodal ; app-visual-pivot ; app-visual-prompt-inds ; app-visual-promptcharm ; app-visual-prompting-inds ; app-visual-survey ; app-visual-understanding ; app-visual-vp3d However, training visual-prompting models poses new challenges despite their practical relevance due to the high diversity and intra-class variation among visual exemplars. Compared to their language-prompted counterparts gdino ; gdino1.5 ; mmgdino ; cp-detr ; ovd-prompt , training recipes for visual prompting detectors remain largely underdeveloped. We have observed that training with supervised fine-tuning (SFT) alone typically struggles under these conditions, exhibiting unstable convergence, limited generalization to out-of-domain data, as well as poor alignment of query predictions with visual prompts (Sec. 4). These findings suggest that vanilla SFT is insufficient for effectively training visual prompting detectors, motivating us to explore fundamentally new training strategies capable of effectively reasoning over diverse visual inputs for robust generalization.
@@
Inspired by the recent breakthroughs in RL-based training frameworks for LRMs bi2024deepseek ; lu2024deepseek ; guo2025deepseek ; liu2024deepseek ; grpo ; ppo ; rlhf ; dpo ; llama2 ; llava ; llama3 ; qwen ; qwen2 , which efficaciously exploit large-scale noisy training data, we aim to similarly unlock the potential of reasoning capabilities within pure vision models, e.g., VFMs. Yet, naïve application of language-based RL methods such as GRPO to vision presents non-trivial challenges. For one thing, GRPO assumes the model behaves as a probabilistic generator, explicitly sampling diverse outputs from learned distributions per input. In contrast, vision models typically produce deterministic structured predictions, making it nontrivial to optimize over a sampled output space. For another, GRPO’s KL-regularization that stabilizes training via constraining token-level output distributions in language models can not be easily translated to structured visual predictions due to fundamental differences in language and vision formulation.
@@ 2 Related Work
Vision foundation models (VFMs) clip ; dinov ; dinox ; florence ; gdino ; gdino1.5 ; glip ; internvl ; sam ; sam2 ; swin ; trex ; trex2 ; vlmr1 ; yoloworld have enabled significant progress across a wide spectrum of visual tasks by learning general-purpose image representations from large-scale datasets. Among them, DETR detr and its derivatives dndetr ; dabdetr ; dino ; dinov ; gdino ; ovdetr-cm ; dinox ; gdino1.5 ; mmgdino formulate object detection as a set prediction problem using transformer-based architectures, offering open-set capabilities and strong performance on dense prediction tasks. Grounding DINO gdino ; mmgdino , in particular, extends DETR to open-vocabulary detection by incorporating vision-language alignment, enabling phrase-grounded object detection via language prompts.
@@
Open-vocabulary detection ovrcnn (OVD) aims to recognize categories beyond the supervised training set by leveraging external knowledge sources such as pretrained text encoders or image-text pairs. Most existing OVD methods ovd-distill ; ovd-prompt ; ovdetr-cm ; ovrcnn ; ovseg-diff ; gdino focus on language prompts to bridge the category gap, leaving the space of visual prompting—using visual exemplars instead of text—largely underexplored. Recent works cp-detr ; dinov ; trex ; trex2 ; dinox have explored using reference images or bounding boxes to ground object-level semantics. However, these methods often rely on inference-time conditioning and lack a robust training paradigm for learning from high-variance visual prompts.
@@
Reinforcement learning rlhf ; ppo ; dpo ; grpo ; bi2024deepseek has played a central role in fine-tuning large language models (LLMs)llama2 ; llama3 ; bert ; qwen ; qwen2 ; bi2024deepseek ; guo2025deepseek via methods such as Reinforcement Learning from Human Feedback (RLHF) rlhf and Group Relative Policy Optimization (GRPO)grpo ; guo2025deepseek . These approaches enable models to align better with diverse, weakly-supervised, or ambiguous objectives. However, the application of reinforcement-style training to vision foundation models, especially for dense prediction tasks like object detection, remains underexplored. Our work bridges this gap by adapting GRPO principles to query-level learning in transformer-based object detectors through our proposed Group Relative Query Optimization.
@@ 3 Methodology
While language-prompted object detection has received growing attention in the vision-language community, the training strategy for visual prompt-based detection remains underexplored. To address the challenge of high-variance visual prompts and unlock the potential of prompt-guided detector, in this work, we introduce a novel training paradigm for the visual prompting object detection, Group Relative Query Optimization (GRQO). Built on top of the Grounding-DINO (G-DINO) gdino framework (§3.1), our approach integrates a visual prompt encoder (§3.2) and the GRQO mechanism (§3.3) to enhance query learning and improve detection robustness.
@@ 3.2 VIS-G-DINO
[…]
@@ 3.3 Group Relative Query Optimization
Visual prompting detection demands that object queries align with highly diverse visual exemplars that share the same semantics. This setting introduces greater intra-class variance than language prompts, requiring the model to both memorize diverse appearances and generalize across unseen variations. Inspired by the generalization ability of GRPO grpo in LLMs community, we propose Group Relative Query Optimization (GRQO, Fig.2)—a novel training paradigm that enhances query quality and learning stability through group-based reward modeling and distributional regularization.
@@
where $\alpha$ and $\beta$ are scalar weights that balance the reward signal and regularization strength. To mitigate the impact of unreliable signals, we set $\alpha$ to zero for the final set of queries exhibiting low objectness probability. GRQO incentivizes both query-level learning and stable objectness modeling.
@@ 4.1 Experiment Setup
We compare GRQO with standard supervised fine-tuning (SFT). All experiments are conducted using the MM-Grounding-DINO mmgdino implementation, which we adapt to support visual prompting. We use visual exemplars (images with user defined bounding boxes) as input prompts to guide the detection in the target image.
@@
We evaluate our approach under two settings: ❶ Zero-Shot. (Out-domain Evaluation). We conduct open-vocabulary detection where we train on Objects365 (O365) o365 and test on COCO cocods , LVIS-minival lvis , ODinW13 glip , and ODinW35 glip . We train the model for 6 epochs as SFT baselines. For GRQO, we first train 1 epoch with SFT to obtain the reference model, then apply GRQO for 5 additional epochs. ❷ Fine-Tune. (In-domain Evaluation). We fine-tune on the COCO training split for 12 epochs and evaluate on the COCO validation set. For GRQO, we apply both pretrained weights of SFT and GRQO for evaluation.
@@ 4.2 Main Results
GDINO-T gdino O365,G,C 30 48.4 28.8 18.8 24.2 34.7 51.4 22.7 58.1
MM-GDINO-T mmgdino O365,G 30 50.4 35.7 28.1 30.2 42.0 45.3 20.2 58.2
MM-GDINO-L mmgdino O365v2,O,G 42 53.0 - - - - 75.3 35.2 -